OpenStudy (anonymous):

If cos x = .67, what is sin x?

OpenStudy (anonymous):

Is this with or without a scientific calculator?

OpenStudy (anonymous):

use sinx =sqrt(1-cos^2x) =sqrt(1-(.67*.67))

OpenStudy (anonymous):

matricked's answer assumes a lot of trigonometry knowledge. How much trigonometry have you done?

hero (hero):

\[\cos^2(x) + \sin^2(x) = 1 \\(.67)^2 + sin^2(x) = 1 \\\sin^2(x) = 1 - (.67)^2 \\\sin(x) = \sqrt{1 - (.67)^2} \]